OpenStudy (anonymous):

Determine if the sequence 1, 2, 3, 5, 8, 13... is arithmetic, geometric, or niether. A. Arithmetic B. Geometric C. Niether

OpenStudy (anonymous):

Assume its an arithmetic ,then d= t2-t1 con.. Let's see , 2-1=1 , 3-2=1 ,5-3=2 So no its not arithmetic . Assume its geometric ,thend t2/t1 =r con .. let's see ,2/1=2 , 3/2=1.5 , 5/3 =1..6 So no its not geometric .